OpenStudy (anonymous):

Which expressions have a value of 42? Choose all answers that are correct. 105 – 32 • (5 + 2) 80 ÷ 2 • (4 + 6) + 2 3 • [6 + 2 • (7 – 3)] [102 – (10 – 2) • 9] + 12

OpenStudy (anonymous):

Do the work. And which one do you think it is..

OpenStudy (anonymous):

105 - 32 * (5 + 2) = 105 - 32 * 8 = 105 - 256 = - 151 (not this one) 80/2 * (4 + 6) + 2 = 40 * 10 + 2 = 400 + 2 = 402 (nope) 3 * [ 6 + 2 * (7 - 3) ] = 3 * [ 6 + 2*4] = 3 * [ 6 + 8 ] = 3 * 14 = 42 (BINGO) [ 102 - (10 - 2) * 9 ] + 12 = [ 102 - 8 * 9 ] + 12 = [ 102 - 72 ] + 12 = 30 + 12 = 42 (BINGO)

OpenStudy (anonymous):

THANK YOU SO MUCH!! :) and @juliephillips really? dont cmment on my stuff with that. I put it on here cuz i needed help and didnt know.

OpenStudy (anonymous):

Sorry..

OpenStudy (anonymous):

its ok. im not mad i just dont understand why people do that lol